# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ III-E Optimization Solution for Weighted Sum-Power Minimization
To solve the problem in (37), we divide it into two sub-problems and solve them alternately through a series of iterations. The first sub-problem is to solve for ${\bf q}_{1}$ and ${\bf F}_{1}$ given that ${\bf F}_{2}\in\mathcal{F}_{2}$ is fixed, and the second sub-problem is to solve for ${\bf F}_{2}\in\mathcal{F}_{2}$ assuming ${\bf q}_{1}$ and ${\bf F}_{1}$ are fixed.
@@ III-E1 First sub-problem for obtaining ${\bf q}_{1}$ and ${\bf F}_{1}$
We assume a fixed value for ${\bf F}_{2}\in\mathcal{F}_{2}$. We first show that $\mathbb{E}[\|{\bf x}_{2}\|^{2}]$ is upper bounded by the sum of the scaled version of $\mathbb{E}[\|{\bf x}_{1}\|^{2}]$ and some constant terms as follows:
$\mathbb{E}[\|{\bf x}_{2}\|^{2}]\overset{(i)}{=}\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}{\bf Q}_{1}^{1/2}{\bf q}_{1}\|^{2}+\|{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2}$
$\hskip 14.22636pt~{}~{}~{}~{}~{}~{}~{}\leq\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+f_{2,\max}^{2}\big{(}\|{\bf Q}_{1}^{1/2}{\bf q}_{1}\|^{2}+\|{\bf F}_{1}{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2}\big{)}$
$\hskip 14.22636pt~{}~{}~{}~{}~{}~{}~{}=\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+f_{2,\max}^{2}\mathbb{E}[\|{\bf x}_{1}\|^{2}],$ (38)
where $f_{2,{\rm max}}^{2}=\underset{i=2,...,N-1}{\max}f_{2,i}^{2}$. To obtain the equality $(i)$ in (38), we use the fact that $\textrm{tr}({\bf F}_{2}{\bf F}_{1}{\bf F}_{2}{\bf F}_{2}^{\top})=0$ for ${\bf F}_{2}\in\mathcal{F}_{2}$. To derive the inequality in (38), we exploit the property that $\|{\bf F}_{2}{\bf y}\|^{2}\leq f_{2,{\rm max}}^{2}\|{\bf y}\|^{2}$ for any vector ${\bf y}\in\mathbb{R}^{N\times 1}$ and $\|{\bf F}_{2}{\bf Y}\|_{F}^{2}\leq f_{2,{\rm max}}^{2}\|{\bf Y}\|_{F}^{2}$ for any matrix ${\bf Y}\in\mathbb{R}^{N\times N}$ when ${\bf F}_{2}\in\mathcal{F}_{2}$. Accordingly, we obtain an upper bound on the objective function in (37) as
${(1-\alpha)(\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2})}+(\alpha+f_{2,\max}^{2}(1-\alpha))\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}.$ (39)
In the first sub-problem, instead of solving (37) directly, we aim to minimize the upper bound of the objective function in (39). Since the other terms in (39) are constants except for $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$, the first sub-problem is then reduced to
$(\bm{\mathcal{{P}}_{1}}):\hskip 5.69046pt$ $\underset{{\bf q}_{1},{\bf F}_{1}}{\text{minimize}}$ $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$
@@
We will solve $\bm{\mathcal{{P}}_{1}}$ via (i) obtaining the optimal solution form of ${\bf F}_{1}$ in terms of ${\bf q}_{1}$, and then (ii) plugging the optimal solution form of ${\bf F}_{1}$ in $\mathbb{E}[\|{\bf x}_{1}\|^{2}]$ and solving the problem for ${\bf q}_{1}$.
@@
where $i\in\{0,...,N-1\}$. With the defined vectors $\{{\bf f}_{1,i}\}$ and $\{{\bf h}_{i}\}$, we can rewrite $\mathbb{E}[\|{\bf x}_{1}\|^{2}]$ in (35) as
$\mathbb{E}[\|{\bf x}_{1}\|^{2}]=\sum_{i=1}^{N-1}\Phi_{i}({\bf f}_{1,i})+\sigma_{1}^{2}\big{(}q_{1,N-1}^{2}+q_{1,N}^{2}\big{)},$ (42)
where $\Phi_{1}({\bf f}_{1,1})\triangleq({\bf h}_{1}^{\top}{\bf f}_{1,1})^{2}\sigma_{2}^{2}+{\bf f}_{1,1}^{\top}{\bf f}_{1,1}\sigma_{2}^{2}$ and $\Phi_{i}({\bf f}_{1,i})\triangleq\big{(}q_{1,i-1}+f_{2,i}{\bf h}_{i}^{\top}{\bf f}_{1,i}\big{)}^{2}\sigma_{1}^{2}+({\bf h}_{i}^{\top}{\bf f}_{1,i})^{2}\sigma_{2}^{2}+{\bf f}_{1,i}^{\top}{\bf f}_{1,i}(f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2})$, $i\in\{2,\cdots,N-1\}$. The detailed derivation of $\Phi_{i}({\bf f}_{1,i})$, $i\in\{1,\cdots,N-1\}$, is presented in Appendix H.
Using (42), our problem of interest (i.e., $\underset{{\bf F}_{1}}{\text{min}}~{}\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]})$ can be decomposed into $N-1$ independent sub-problems each in the form of $\underset{{\bf f}_{1,i}}{\text{min}}~{}\Phi_{i}({\bf f}_{1,i})$, $i\in\{1,...,N-1\}$. Since each independent sub-problem is convex with respect to ${\bf f}_{1,i}$, we find ${\bf f}_{1,i}$ optimally by solving $\frac{\partial\Phi_{i}({\bf f}_{1,i})}{\partial{\bf f}_{1,i}}={\bf 0}^{\top}$. Obviously, we have ${\bf f}_{1,1}={\bf 0}$. Also, for $i\in\{2,...,N-1\}$, we need to solve
@@
For a detailed derivation of (44) from (43), refer to Appendix I.
@@
$\mathbb{E}[\|{\bf x}_{1}\|^{2}]$ $=\sum_{i=2}^{N-1}\bigg{[}\bigg{(}q_{1,i-1}-\frac{f_{2,i}^{2}\sigma_{1}^{2}}{f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2}}\frac{q_{1,i-1}\|{\bf h}_{i}\|^{2}}{1+\|{\bf h}_{i}\|^{2}}\bigg{)}^{2}\sigma_{1}^{2}+\bigg{(}\frac{f_{2,i}\sigma_{1}^{2}}{f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2}}\frac{q_{1,i-1}\|{\bf h}_{i}\|^{2}}{1+\|{\bf h}_{i}\|^{2}}\bigg{)}^{2}\sigma_{2}^{2}$
$\hskip 85.35826pt+\bigg{(}\frac{f_{2,i}\sigma_{1}^{2}}{f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2}}\frac{q_{1,i-1}}{1+\|{\bf h}_{i}\|^{2}}\bigg{)}^{2}\|{\bf h}_{i}\|^{2}(f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2})\bigg{]}+\big{(}q_{1,N-1}^{2}+q_{1,N}^{2}\big{)}\sigma_{1}^{2}$
@@
Then, $\bm{\mathcal{{P}}_{1}}$ can be reduced to the following optimization problem:
@@ III-E2 Second sub-problem for obtaining ${\bf F}_{2}$
$(\bm{\mathcal{{P}}_{2}}):\hskip 5.69046pt\underset{{\bf F}_{2}\in\mathcal{F}_{2}}{\text{minimize}}$ $\alpha\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}+(1-\alpha)\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}.$ (49)
We aim to minimize the objective of $\bm{\mathcal{{P}}_{2}}$ for each $f_{2,i}$, $i\in\{2,...,N-1\}$ by setting the derivative with respect to $f_{2,i}$ equal to zero. Our methodology would yield a sub-optimal solution given the non-convexity of the problem $\bm{\mathcal{{P}}_{2}}$. We then obtain
$\alpha\frac{\partial\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}}{\partial f_{2,i}}+(1-\alpha)\frac{\partial\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}}{\partial f_{2,i}}=2\alpha\sigma_{1}^{2}q_{1,i-1}{\bf h}_{i}^{\top}{\bf f}_{1,i}+c_{i}f_{2,i},$ (50)
@@
The details of derivations used for obtaining (50) are presented in Appendix J. By setting the right-hand side of the equality in (50) to zero, we obtain the solution for $f_{2,i}$ as
@@
The pseudo-code of our iterative method to solve the overall optimization problem in (29) is summarized in Algorithm 1. We solve the two sub-problems alternatively through a series of outer iterations denoted in lines 5-17. In the inner iterations, lines 11-13, we solve the second sub-problem.
@@
3:Obtain the optimal solutions, ${\bf g}_{2}=[0,0,...,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$ and ${f}_{2,N}=0$, from Propositions 1 and 2.
@@
9:    Obtain ${\bf g}_{1}={\bf Q}_{1}^{1/2}{\bf q}_{1}$ where ${\bf Q}_{1}$ is given in (20).

### v2

@@ III-E Optimization Solution for Weighted Sum-Power Minimization
To solve the problem in (55), we divide it into two sub-problems and solve them alternately through a series of iterations. [^3] Although this iterative optimization process does not guarantee convergence, it simplifies (55) into two more tractable sub-problems. We demonstrate the convergence empirically through numerical experiments in Sec. VI-J. The first sub-problem is to solve for ${\bf q}_{1}$ and ${\bf F}_{1}$ given that ${\bf F}_{2}\in\mathcal{F}_{2}$ is fixed, and the second sub-problem is to solve for ${\bf F}_{2}\in\mathcal{F}_{2}$ assuming ${\bf q}_{1}$ and ${\bf F}_{1}$ are fixed.
@@ III-E1 First sub-problem for obtaining ${\bf q}_{1}$ and ${\bf F}_{1}$
We assume a fixed value for ${\bf F}_{2}\in\mathcal{F}_{2}$. We first show that $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ is upper bounded by the sum of the scaled version of $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ and some constant terms as follows:
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}[\|{\bf x}_{2}\|^{2}]\overset{(a)}{=}\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}{\bf Q}_{1,\text{sqrt}}{\bf q}_{1}\|^{2}+\|{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2}$
$\hskip 14.22636pt~{}~{}~{}~{}~{}~{}~{}\leq\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+f_{2,\max}^{2}\big{(}\|{\bf Q}_{1,\text{sqrt}}{\bf q}_{1}\|^{2}+\|{\bf F}_{1}{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2}\big{)}$
$\hskip 14.22636pt~{}~{}~{}~{}~{}~{}~{}=\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+f_{2,\max}^{2}\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}[\|{\bf x}_{1}\|^{2}],$ (56)
where $f_{2,{\rm max}}^{2}=\underset{i=2,...,N-1}{\max}f_{2,i}^{2}$. To obtain the equality $(a)$ in (56), we use the fact that $\textrm{tr}({\bf F}_{2}{\bf F}_{1}{\bf F}_{2}{\bf F}_{2}^{\top})=0$ for ${\bf F}_{2}\in\mathcal{F}_{2}$. To derive the inequality in (56), we exploit the property that $\|{\bf F}_{2}{\bf y}\|^{2}\leq f_{2,{\rm max}}^{2}\|{\bf y}\|^{2}$ for any vector ${\bf y}\in\mathbb{R}^{N\times 1}$ and $\|{\bf F}_{2}{\bf Y}\|_{F}^{2}\leq f_{2,{\rm max}}^{2}\|{\bf Y}\|_{F}^{2}$ for any matrix ${\bf Y}\in\mathbb{R}^{N\times N}$ when ${\bf F}_{2}\in\mathcal{F}_{2}$. Accordingly, we obtain an upper bound on the objective function in (55) as
${(1-\alpha)(\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2})}+(\alpha+f_{2,\max}^{2}(1-\alpha))\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}.$ (57)
In the first sub-problem, instead of solving (55) directly, we aim to minimize the upper bound of the objective function in (57). [^4] Although using the upper bound as a surrogate objective may result in sub-optimal solutions, we employ it for tractability in optimization. Since the other terms in (57) are constants except for $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$, the first sub-problem is then reduced to
$({\mathbf{m}}{\mathcal{{P}}_{1}}):\hskip 5.69046pt$ $\underset{{\bf q}_{1},{\bf F}_{1}}{\text{minimize}}$ $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$
@@
We will solve ${\mathbf{m}}{\mathcal{{P}}_{1}}$ via (i) obtaining the optimal solution form of ${\bf F}_{1}$ in terms of ${\bf q}_{1}$, and then (ii) plugging the optimal solution form of ${\bf F}_{1}$ in $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ and solving the problem for ${\bf q}_{1}$.
@@
where $i\in\{0,...,N-1\}$. With the defined vectors $\{{\bf f}_{1,i}\}$ and $\{{\bf h}_{i}\}$, we can rewrite $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ in (53) as
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}=\sum_{i=1}^{N-1}\Phi_{i}({\bf f}_{1,i})+\sigma_{1}^{2}\big{(}q_{1,N-1}^{2}+q_{1,N}^{2}\big{)},$ (60)
where $\Phi_{1}({\bf f}_{1,1})\triangleq({\bf h}_{1}^{\top}{\bf f}_{1,1})^{2}\sigma_{2}^{2}+{\bf f}_{1,1}^{\top}{\bf f}_{1,1}\sigma_{2}^{2}$ and $\Phi_{i}({\bf f}_{1,i})\triangleq\big{(}q_{1,i-1}+f_{2,i}{\bf h}_{i}^{\top}{\bf f}_{1,i}\big{)}^{2}\sigma_{1}^{2}+({\bf h}_{i}^{\top}{\bf f}_{1,i})^{2}\sigma_{2}^{2}+{\bf f}_{1,i}^{\top}{\bf f}_{1,i}(f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2})$, $i\in\{2,\cdots,N-1\}$. The detailed derivation of $\Phi_{i}({\bf f}_{1,i})$, $i\in\{1,\cdots,N-1\}$, is presented in Appendix M.
Using (60), our problem of interest (i.e., $\underset{{\bf F}_{1}}{\text{min}}~{}\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]})$ can be decomposed into $N-1$ independent sub-problems each in the form of $\underset{{\bf f}_{1,i}}{\text{min}}~{}\Phi_{i}({\bf f}_{1,i})$, $i\in\{1,...,N-1\}$. Since each independent sub-problem is convex with respect to ${\bf f}_{1,i}$, we find ${\bf f}_{1,i}$ optimally by solving $\frac{\partial\Phi_{i}({\bf f}_{1,i})}{\partial{\bf f}_{1,i}}={\bf 0}^{\top}$. Obviously, we have ${\bf f}_{1,1}={\bf 0}$. Also, for $i\in\{2,...,N-1\}$, we need to solve
@@
For a detailed derivation of (62) from (61), refer to Appendix N.
@@
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ $=\sum_{i=2}^{N-1}\bigg{[}\bigg{(}q_{1,i-1}-\frac{f_{2,i}^{2}\sigma_{1}^{2}}{f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2}}\frac{q_{1,i-1}\|{\bf h}_{i}\|^{2}}{1+\|{\bf h}_{i}\|^{2}}\bigg{)}^{2}\sigma_{1}^{2}+\bigg{(}\frac{f_{2,i}\sigma_{1}^{2}}{f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2}}\frac{q_{1,i-1}\|{\bf h}_{i}\|^{2}}{1+\|{\bf h}_{i}\|^{2}}\bigg{)}^{2}\sigma_{2}^{2}$
$\hskip 28.45274pt+\bigg{(}\frac{f_{2,i}\sigma_{1}^{2}}{f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2}}\frac{q_{1,i-1}}{1+\|{\bf h}_{i}\|^{2}}\bigg{)}^{2}\|{\bf h}_{i}\|^{2}(f_{2,i}^{2}\sigma_{1}^{2}+\sigma_{2}^{2})\bigg{]}+\big{(}q_{1,N-1}^{2}+q_{1,N}^{2}\big{)}\sigma_{1}^{2}$
@@
Then, ${\mathbf{m}}{\mathcal{{P}}_{1}}$ can be reduced to the following optimization problem:
@@ III-E2 Second sub-problem for obtaining ${\bf F}_{2}$
$({\mathbf{m}}{\mathcal{{P}}_{2}}):\hskip 5.69046pt\underset{{\bf F}_{2}\in\mathcal{F}_{2}}{\text{minimize}}$ $\alpha\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}+(1-\alpha)\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}.$ (67)
We aim to minimize the objective of ${\mathbf{m}}{\mathcal{{P}}_{2}}$ for each $f_{2,i}$, $i\in\{2,...,N-1\}$ by setting the derivative with respect to $f_{2,i}$ equal to zero. Our methodology would yield a sub-optimal solution given the non-convexity of the problem ${\mathbf{m}}{\mathcal{{P}}_{2}}$. We then obtain
$\alpha\frac{\partial\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}}{\partial f_{2,i}}+(1-\alpha)\frac{\partial\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}}{\partial f_{2,i}}=2\alpha\sigma_{1}^{2}q_{1,i-1}{\bf h}_{i}^{\top}{\bf f}_{1,i}+c_{i}f_{2,i},$ (68)
@@
The details of derivations used for obtaining (68) are presented in Appendix O. By setting the right-hand side of the equality in (68) to zero, we obtain the solution for $f_{2,i}$ as
@@
The pseudo-code of our iterative method to solve the overall optimization problem in (46) is summarized in Algorithm 1. We first set ${\bf g}_{2}=[0,0,...,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$ and ${f}_{2,N}=0$, from Propositions 1 and 3, and initialize $\{f_{2,i}\}_{i=2}^{N-1}$ randomly (in lines 3-4). Then, we solve the two sub-problems, discussed in Secs. III-E1 and  III-E2, alternatively through a series of outer iterations until the objective function value in (46) converges (lines 5-18). For the first sub-problem (lines 6-9), we first solve the problem in (66) with respect to ${\bf x}=[x_{1},...,x_{N}]^{\top}$ and obtain $q_{1,i}=\sqrt{x_{i}}$, $i\in\{1,...,N\}$. Then, we obtain the columns of ${\bf F}_{1}$, $\{{\bf f}_{1,i}\}_{i=1}^{N-1}$, from (62), using the values of $\{q_{1,i}\}_{i=1}^{N-1}$, and finally obtain ${\bf g}_{1}={\bf Q}_{1,\text{sqrt}}{\bf q}_{1}$ where ${\bf Q}_{1}$ is given in (31). For the second sub-problem (lines 10-14), we obtain $f_{2,i}$ sequentially for $i\in\{2,...,N-1\}$ by (69) and conduct several inner iterations in obtaining $\{f_{2,i}\}_{i=2}^{N-1}$ until the objective function value in (46) converges.
@@
3:Obtain the optimal solutions, ${\bf g}_{2}=[0,0,...,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$ and ${f}_{2,N}=0$, from Propositions 1 and 3.
@@
9:    Obtain ${\bf g}_{1}={\bf Q}_{1,\text{sqrt}}{\bf q}_{1}$ where ${\bf Q}_{1}$ is given in (31).